-For three days in a row, volunteers picked up trash on a particular beech. They picked up a total of 921 pounds. On the second day they collected 39 pounds more than on the first day. On the third day, they picked up 18 pounds less then on the first day. How many pounds of trash sis they pick up on the first day?
Answer by queenofit(120)

First you know that you set the following according to your problem
x= first day
x+39= second day
x-18= third day
when you add them all together you get a total of 921 pounds. so you must add the three above set = to 921 and solve for x.
x+x+39+x-18=921
3x=921-39+18
3x=900
x=300
you could find the number of pounds for day 2 and day 3 as well. All you would have to do is put 300 into the equation for the day you needed such as
x+39=second day 300+39=339
x-18 third day 300-18=282